# Neutrinoless Double Beta Decay in Light of JUNO First Data

### v1

@@ 2 Effective Majorana Mass for $0\nu 2\beta$ Decay
Translating the measurement of the half-life time $T^{0\nu}_{1/2}$ to the effective mass $\langle m_{ee}\rangle$ suffers from the uncertainty in the nuclear matrix elements (NME) $M^{0\nu}$ in Eq. (1.1). These nuclear physics uncertainties surpass those from neutrino parameter determinations. Following the framework established in [GE1], the Fig. 3 therein is updated as Fig. 3 with the current NME calculations extracted from [Agostini:2022zub]. Fig. 3 illustrates the current experimental limits on $\langle m_{ee}\rangle$ versus $T_{1/2}^{0\nu}$ for six isotopes: 48Ca, 76Ge, 82Se, 100Mo, 130Te, and 136Xe. The light blue bands indicate the nuclear uncertainties using the combined NME range of available calculations for each isotope (see Table I of [Agostini:2022zub]). Even with fixed value for the half-life time $T^{0\nu}_{1/2}$, there is no single fixed value of the effective mass $\langle m_{ee}\rangle$ to be extracted.
@@
The current experimental sensitivities on the half-life time $T^{0\nu}_{1/2}$ at 90% C.L. are represented by the solid vertical lines in Fig. 3. Regions to the left-hand side of the vertical line have been excluded. Among various measurements, the most stringent constraints come from KamLAND-Zen (136Xe) with $T_{1/2}^{0\nu}>2.0\times 10^{26}$ yr [KamLAND-Zen:2024eml, KamLAND-Zen:2022tow, KamLANDZen:2016pfg] and GERDA (76Ge) with $T_{1/2}^{0\nu}>1.2\times 10^{26}$ yr [GERDA:2017foj]. Other competitive bounds include MAJORANA DEMONSTRATOR (76Ge, $T_{1/2}^{0\nu}>3.9\times 10^{25}$ yr) [Majorana:2022udl], EXO-200 (136Xe, $T_{1/2}^{0\nu}>2.4\times 10^{25}$ yr) [EXO200:2019rkq], CUORE (130Te, $T_{1/2}^{0\nu}>4.8\times 10^{25}$ yr) [CUORE:2022xwt], CUPID-Mo (100Mo, $T_{1/2}^{0\nu}>9.7\times 10^{24}$ yr) [CUPID:2022okx], and CUPID-0 (82Se, $T_{1/2}^{0\nu}>4.4\times 10^{24}$ yr) [CUPID:2022okx].
@@
The effort for next-generation experiments promises a qualitative leap in sensitivity, represented by the dashed vertical lines in Fig. 3, with the flagship experiments nEXO (136Xe) [nEXO1, nEXO2] and LEGEND-1000 (76Ge) [LEGEND]. Both are projected to reach an unprecedented half-life sensitivity of $T_{1/2}^{0\nu}\sim 10^{28}$ yr. As illustrated by the intersection of their dashed vertical lines with the light blue NME uncertainty band, the corresponding $\langle m_{ee}\rangle$ sensitivity ranges are $(6\sim 27)$ meV for nEXO and $(9\sim 21)$ meV for LEGEND-1000. In addition, the CUPID experiment [CUPID1, CUPID2] targeting 100Mo and 82Se aims for $T_{1/2}^{0\nu}\sim 10^{27}$ yr which can translate to $\langle m_{ee}\rangle=(12\sim 34)$ meV. We also include in Fig. 3 the projected sensitivities of DARWIN [DARWIN2], KamLAND-Zen [KamLANDZen6], NEXT-HD [NEXT2], SNO+ [SNOplus2], AMoRE-II [AMoRE], SuperNEMO [SuperNEMO1], and LEGEND-200 [LEGEND].
@@ 4.1 Cosmological Measurement of Neutrino Mass Scale and Preference for NO
The cosmological constraints on the neutrino mass sum $\sum_{i}m_{i}$ provide complementary information to laboratory searches. Especially, the DESI BAO measurements with CMB data gives the most stringent current bound, $\sum_{i}m_{i}<0.082$ eV at $95\%$ C.L. [DESI2024b, DESI2024]. For reference, the combination of CMB data from Planck, the non-DESI Baryon Acoustic Oscillations (BAO) observation, Cosmic Chronometers (CC), and Pantheon, yields $\sum_{i}m_{i}<0.12$ eV at $95\%$ C.L. [Planck:2018vyg]. These bounds are indicated in the left panel of Fig. 4 with vertical lines. The current cosmological observations have already went beyond the IO regime to probe its NO counterpart. The future cosmological surveys, such as EUCLID, are expected to reach the sensitivity of $\sum_{i}m_{i}<0.056$ eV [EUCLID2020, EUCLID2022] which would directly test the global lower limit of the neutrino mass sum.
@@
The kinematic constraint $m_{\beta}<0.45$ eV at $90\%$ C.L. [KATRIN2022, KATRIN2024] on the beta decay effective mass $m_{\beta}\equiv\sqrt{\sum_{i}|U_{ei}|^{2}m_{i}^{2}}$ from the KATRIN experiment provides a direct and model-independent measurement of the neutrino mass scale. Although the beta decay constraint is not as strong as the cosmological observation, we also show them in the right panel of Fig. 4 for comparison. The current laboratory measurements probe the cosmologically allowed mass range, but still remain about an order of magnitude above the sensitivity required to test the IO scenario. The KATRIN experiment is expected to reach its design sensitivity $m_{\beta}\sim 0.2$ eV at $90\%$ C.L. after several years of data taking [KATRINFuture]. The next-generation experiments such as PROJECT 8 are projected to reach the sensitivity of $m_{\beta}\sim 0.04$ eV [Project8, Project8Roadmap], which would probe the majority of the IO regime.
@@
The behavior of $\langle m_{ee}\rangle$ in the complex plane reveals the fundamental difference between the two mass orderings. For NO, the mass hierarchy $m_{1}<m_{2}\ll m_{3}$ allows these three vectors to form a closed triangle that can reach complete cancellation, $\langle m_{ee}\rangle=0$, for specific combinations of the two Majorana CP phases [Vissani:1999tu, Xing:2014yka, GE2]. In contrast, for IO with $m_{3}\ll m_{1}<m_{2}$, the relative magnitudes of the vector sides prevent such closure, making $\langle m_{ee}\rangle=0$ geometrically impossible regardless of the Majorana phases. Such fundamental difference manifests topologically in the complex $\langle m_{ee}\rangle$ plane for NO.
@@
Across the three panels of Fig. 5, the size of the circles differ considerably. The major factor is the allowed range of the neutrino mass scale. For comparison, the left panel shows the results obtained with even distribution (flat prior) of the lightest mass in the range of $m_{1}\in[1,50]$ meV. Once the cosmological constraints are imposed in the center and right panels, the allowed circle regions shrink. The effect is not that significant for the case with only pre-DESI constraints since the limit $\sum_{i}m_{i}<0.12$ eV 95% C.L. [Planck:2018vyg] translates to $m_{1}\lesssim 30$ meV which is roughly the size of circle. To make the connection more direct, the corresponding upper limit for the NO effective mass in Eq. (2.8) with a nonzero $m_{1}$ is,
@@
However, once the DESI constraint $\sum m_{i}<0.082$ eV at 95% C.L.[DESI2024b, DESI2024] is imposed, the allowed circle significantly shrinks to $\langle m_{ee}\rangle^{\rm NO}_{\rm max}(m_{1})\approx 16$ meV with $m_{1}\lesssim 14$ meV.
@@ 4.2 Determination of Two Majorana CP Phases
If the cancellation among the three complex vectors $\overrightarrow{L_{i}}$ really happens, a Majorana triangle can form. Then it is possible to determine the two Majorana CP phases simultaneously as function of the three vector lengths [GE2],
@@
A paradox here is that if the $0\nu 2\beta$ decay is observed, one physical degree of freedom cannot be obversed. But if the effective mass $\langle m_{ee}\rangle$ vanishes to make the $0\nu 2\beta$ decay invisible, both physical Majorana CP phases can be simultaneously determined. The Majorana nature can be probed with other ways, such as the direct detection of cosmic neutrino backgrounds with tritium [Ringwald:2009bg, Vogel:2015vfa, Shergold:2021evs] or the atomic radiative emission of neutrino pairs [Dinh:2012qb, Fukumi:2012rn].
@@
Fig. 7 illustrates this possibility by showing the two Majorana CP phases $\delta_{\rm M1}$ and $\delta_{\rm M3}$ determined by the Majorana triangle ($\langle m_{ee}\rangle=0$) for NO with different values of the lightest neutrino mass $m_{1}=3$, 4, 5, 6 meV. Each colored region represents one allowed solution: blue for 3 meV, orange for 4 meV, green for 5 meV, and red for 6 meV. These solutions are obtained by varying the oscillation parameters within $\Delta\chi^{2}<9$ and solving Eq. (4.2) for each parameter combination. Three uncertainty schemes are displayed: the previous uncertainties used in the Fig. 7 of [GE2] (dotted contours), current NuFIT 6.0 bounds [NuFIT6] (dashed contours), and the combination of NuFIT with the JUNO first measurement [JUNO1st] (thick solid contours). The progressive narrowing of these allowed regions, particularly evident when comparing the dotted with the solid contours, directly demonstrates how the reduced uncertainties of oscillation parameters translate into more precise predictions for the Majorana phases solutions.
@@
Fig. 7: The two Majorana CP phases $(\delta_{\rm M1},\delta_{\rm M3})$ determined by the Majorana triangle ($\langle m_{ee}\rangle=0$) for NO with $m_{1}=3$ meV (blue), 4 meV (orange), 5 meV (green), and 6 meV (red). Three uncertainty schemes of the oscillation parameters with $\Delta\chi^{2}<9$ are shown. The dotted contours represent results from the Fig. 7 of [GE2], dashed contours show the current NuFIT 6.0 bounds, while the thick solid contours incorporate the latest JUNO measurement.
@@
The comparison in Fig. 7 clearly demonstrates how the reactor experiments can help mitigating the uncertainty in $0\nu 2\beta$ decay. The factor of $\sqrt{2}$ reduction in the $\theta_{s}$ uncertainty at JUNO narrows the allowed region of $\delta_{\rm M1}$ and $\delta_{\rm M3}$ for each value of $m_{1}$. This reduction is particularly pronounced because $\theta_{s}$ directly affects the relative magnitudes of $L_{1}$ and $L_{2}$ through the prefactors $c_{s}^{2}$ and $s_{s}^{2}$. The narrower bands in Fig. 7 translate directly into more precise predictions. If future $0\nu\beta\beta$ experiments do not observe a signal despite reaching sensitivities $|\langle m_{ee}\rangle|\lesssim 1$ meV, JUNO’s improved oscillation parameters would enable a much sharper determination of the Majorana CP phases under the assumption of complete cancellation.

### v2

@@ 2 Effective Majorana Mass for $0\nu 2\beta$ Decay
Translating the measurement of the half-life time $T^{0\nu}_{1/2}$ to the effective mass $\langle m_{ee}\rangle$ suffers from the uncertainty in the nuclear matrix elements (NME) $M^{0\nu}$ in Eq. (1.1). These nuclear physics uncertainties surpass those from neutrino parameter determinations. Following the framework established in [57], the Fig. 3 therein is updated as Fig. 3 with the current NME calculations extracted from [24]. Fig. 3 illustrates the current experimental limits on $\langle m_{ee}\rangle$ versus $T_{1/2}^{0\nu}$ for six isotopes: 48Ca, 76Ge, 82Se, 100Mo, 130Te, and 136Xe. The light blue bands indicate the nuclear uncertainties using the combined NME range of available calculations for each isotope (see Table I of [24]). Even with fixed value for the half-life time $T^{0\nu}_{1/2}$, there is no single fixed value of the effective mass $\langle m_{ee}\rangle$ to be extracted.
@@
The current experimental sensitivities on the half-life time $T^{0\nu}_{1/2}$ at 90% C.L. are represented by the solid vertical lines in Fig. 3. Regions to the left-hand side of the vertical line have been excluded. Among various measurements, the most stringent constraints come from KamLAND-Zen (136Xe) with $T_{1/2}^{0\nu}>2.0\times 10^{26}$ yr [77, 78, 85] and GERDA (76Ge) with $T_{1/2}^{0\nu}>1.2\times 10^{26}$ yr [86]. Other competitive bounds include MAJORANA DEMONSTRATOR (76Ge, $T_{1/2}^{0\nu}>3.9\times 10^{25}$ yr) [87], EXO-200 (136Xe, $T_{1/2}^{0\nu}>2.4\times 10^{25}$ yr) [88], PandaX-4T (136Xe, $T_{1/2}^{0\nu}>2.1\times 10^{24}$ yr) [89], CUORE (130Te, $T_{1/2}^{0\nu}>4.8\times 10^{25}$ yr) [90], CUPID-Mo (100Mo, $T_{1/2}^{0\nu}>9.7\times 10^{24}$ yr) [91], and CUPID-0 (82Se, $T_{1/2}^{0\nu}>4.4\times 10^{24}$ yr) [91].
@@
The effort for next-generation experiments promises a qualitative leap in sensitivity, represented by the dashed vertical lines in Fig. 3, with the flagship experiments nEXO (136Xe) [79, 80] and LEGEND-1000 (76Ge) [81]. Both are projected to reach an unprecedented half-life sensitivity of $T_{1/2}^{0\nu}\sim 10^{28}$ yr. As illustrated by the intersection of their dashed vertical lines with the light blue NME uncertainty band, the corresponding $\langle m_{ee}\rangle$ sensitivity ranges are $(6\sim 27)$ meV for nEXO and $(9\sim 21)$ meV for LEGEND-1000. In addition, the CUPID experiment [83, 84] targeting 100Mo and 82Se aims for $T_{1/2}^{0\nu}\sim 10^{27}$ yr which can translate to $\langle m_{ee}\rangle=(12\sim 34)$ meV. We also include in Fig. 3 the projected sensitivities of DARWIN [92], KamLAND-Zen [93], NEXT-HD [94], SNO+ [95], AMoRE-II [96], SuperNEMO [97], and LEGEND-200 [81].
@@ 4.1 Cosmological Measurement of Neutrino Mass Scale and Preference for NO
The cosmological constraints on the neutrino mass sum $\sum_{i}m_{i}$ provide complementary information to laboratory searches. Especially, the DESI BAO measurements with CMB data gives the most stringent current bound, $\sum_{i}m_{i}<0.082$ eV at $95\%$ C.L. [70, 69]. For reference, the combination of CMB data from Planck, the non-DESI Baryon Acoustic Oscillations (BAO) observation, Cosmic Chronometers (CC), and Pantheon, yields $\sum_{i}m_{i}<0.12$ eV at $95\%$ C.L. [71]. These bounds are indicated in the left panel of Fig. 4 with vertical lines. The current cosmological observations have already went beyond the IO regime to probe its NO counterpart. The future cosmological surveys, such as EUCLID, are expected to reach the sensitivity of $\sum_{i}m_{i}<0.056$ eV [98, 99] which would directly test the global lower limit of the neutrino mass sum.
@@
The kinematic constraint $m_{\beta}<0.45$ eV at $90\%$ C.L. [67, 68] on the beta decay effective mass $m_{\beta}\equiv\sqrt{\sum_{i}|U_{ei}|^{2}m_{i}^{2}}$ from the KATRIN experiment provides a direct and model-independent measurement of the neutrino mass scale. Although the beta decay constraint is not as strong as the cosmological observation, we also show them in the right panel of Fig. 4 for comparison. The current laboratory measurements probe the cosmologically allowed mass range, but still remain about an order of magnitude above the sensitivity required to test the IO scenario. The KATRIN experiment is expected to reach its design sensitivity $m_{\beta}\sim 0.2$ eV at $90\%$ C.L. after several years of data taking [110]. The next-generation experiments such as PROJECT 8 are projected to reach the sensitivity of $m_{\beta}\sim 0.04$ eV [111, 112], which would probe the majority of the IO regime.
@@
The behavior of $\langle m_{ee}\rangle$ in the complex plane reveals the fundamental difference between the two mass orderings. For NO, the mass hierarchy $m_{1}<m_{2}\ll m_{3}$ allows these three vectors to form a closed triangle that can reach complete cancellation, $\langle m_{ee}\rangle=0$, for specific combinations of the two Majorana CP phases [73, 74, 58]. In contrast, for IO with $m_{3}\ll m_{1}<m_{2}$, the relative magnitudes of the vector sides prevent such closure, making $\langle m_{ee}\rangle=0$ geometrically impossible regardless of the Majorana phases. Such fundamental difference manifests topologically in the complex $\langle m_{ee}\rangle$ plane for NO.
@@
Across the three panels of Fig. 5, the size of the circles differ considerably. The major factor is the allowed range of the neutrino mass scale. For comparison, the left panel shows the results obtained with even distribution (flat prior) of the lightest mass in the range of $m_{1}\in[1,50]$ meV. Once the cosmological constraints are imposed in the center and right panels, the allowed circle regions shrink. The effect is not that significant for the case with only pre-DESI constraints since the limit $\sum_{i}m_{i}<0.12$ eV 95% C.L. [71] translates to $m_{1}\lesssim 30$ meV which is roughly the size of circle. To make the connection more direct, the corresponding upper limit for the NO effective mass in Eq. (2.8) with a nonzero $m_{1}$ is,
@@
However, once the DESI constraint $\sum m_{i}<0.082$ eV at 95% C.L.[70, 69] is imposed, the allowed circle significantly shrinks to $\langle m_{ee}\rangle^{\rm NO}_{\rm max}(m_{1})\approx 16$ meV with $m_{1}\lesssim 14$ meV.
@@ 4.2 Determination of Two Majorana CP Phases
If the cancellation among the three complex vectors $\overrightarrow{L_{i}}$ really happens, a Majorana triangle can form. Then it is possible to determine the two Majorana CP phases simultaneously as function of the three vector lengths [58],
@@
A paradox here is that if the $0\nu 2\beta$ decay is observed, one physical degree of freedom cannot be observed. But if the effective mass $\langle m_{ee}\rangle$ vanishes to make the $0\nu 2\beta$ decay invisible, both physical Majorana CP phases can be simultaneously determined. The Majorana nature can be probed with other ways, such as the direct detection of cosmic neutrino backgrounds with tritium [113, 114, 23] or the atomic radiative emission of neutrino pairs [115, 116].
Besides [58], determining the Majorana CP phases according to Eq. (4.2) as cosine functions of the Majorana triangle sides has also been explored in [56] and [59]. Both [58] and [59] has projected the uncertainty reduction by the JUNO experiment. In our current study, we would use the first data release from JUNO [66] to make realistic illustration of the improvement that the JUNO experiment can already provide.
@@
Fig. 7 illustrates this possibility by showing the two Majorana CP phases $\delta_{\rm M1}$ and $\delta_{\rm M3}$ determined by the Majorana triangle ($\langle m_{ee}\rangle=0$) for NO with different values of the lightest neutrino mass $m_{1}=3$, 4, 5, 6 meV. Each colored region represents one allowed solution: blue for 3 meV, orange for 4 meV, green for 5 meV, and red for 6 meV. These solutions are obtained by varying the oscillation parameters within $\Delta\chi^{2}<9$ and solving Eq. (4.2) for each parameter combination. Three uncertainty schemes are displayed: the previous uncertainties used in the Fig. 7 of [58] (dotted contours), current NuFIT 6.0 bounds [65] (dashed contours), and the combination of NuFIT with the JUNO first measurement [66] (thick solid contours). The progressive narrowing of these allowed regions, particularly evident when comparing the dotted with the solid contours, directly demonstrates how the reduced uncertainties of oscillation parameters translate into more precise predictions for the Majorana phases solutions.
@@
Fig. 7: The two Majorana CP phases $(\delta_{\rm M1},\delta_{\rm M3})$ determined by the Majorana triangle ($\langle m_{ee}\rangle=0$) for NO with $m_{1}=3$ meV (blue), 4 meV (orange), 5 meV (green), and 6 meV (red). Three uncertainty schemes of the oscillation parameters with $\Delta\chi^{2}<9$ are shown. The dotted contours represent results from the Fig. 7 of [58], dashed contours show the current NuFIT 6.0 bounds, while the thick solid contours incorporate the latest JUNO measurement.
@@
The comparison in Fig. 7 clearly demonstrates how the reactor experiments can help mitigating the uncertainty in $0\nu 2\beta$ decay. The factor of $\sqrt{2}$ reduction in the $\theta_{s}$ uncertainty at JUNO narrows the allowed region of $\delta_{\rm M1}$ and $\delta_{\rm M3}$ for each value of $m_{1}$. This reduction is particularly pronounced because $\theta_{s}$ directly affects the relative magnitudes of $L_{1}$ and $L_{2}$ through the prefactors $c_{s}^{2}$ and $s_{s}^{2}$. The narrower bands in Fig. 7 translate directly into more precise predictions. If future $0\nu\beta\beta$ experiments do not observe a signal despite reaching sensitivities $|\langle m_{ee}\rangle|\lesssim 1$ meV, JUNO’s improved oscillation parameters would enable a much sharper determination of the Majorana CP phases under the assumption of complete cancellation. The effect of the $\langle m_{ee}\rangle$ experimental uncertainties on the Majorana CP phase determination can be found in [58] and [61].